OpenStudy (anonymous):

Express the complex number in trigonometric form. 3

OpenStudy (anonymous):

\[\Large z=x+yi \\ \Large z=3 \] Now the rest is only trigonometry in the xy-Plane.

OpenStudy (anonymous):

it is just another way of writing 3 since \(\cos(0)=1\) and \(\sin(0)=0\) you have \[3\left(\cos(0)+i\sin(0)\right)=3\]

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

Now what if instead of 3 it's like 2i ?

OpenStudy (anonymous):

sorry i didn't know someone else was answering